OpenStudy (anonymous):

Find the equation of the tangent to the curve y = cosπx at the point when x = ½

OpenStudy (anonymous):

You need a point....first do that by plugging in x = ½ to y = cosπx (½,y) where y is whatever you find Afterwards you need a slope....you do that by finding the derivative of cosπx and plugging in x = ½ After that you use y-y1=m(x-x1) where (x1,y1) is the point you got and m is the slope

OpenStudy (anonymous):

Yeah, I've done that. I got: y = cosπ/2 m = -πsinπx y - cosπ/2 = -πsinπ/2(x - ½) 2y - cosπ = -πsinπ(x - ½) 2y - cosπ = -πxsinπ + (πsinπ)/2 4y - 2cosπ = -2πxsinπ + πsinπ 2πxsinπ + 4y - πsinπ - 2cosπ = 0 I think I'm wrong, I feel i did it wrong... help haha.

OpenStudy (anonymous):

Looking at the first part it seems like you did it right but I'll get someone else to check....I think it looks right

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

i guess you were right